40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jan 17 15:07:28 2015, manasa, Update, General, Beat note frequency discrepancy 
    Reply  Mon Jan 19 11:03:32 2015, manasa, Update, General, RF connections to adder 
    Reply  Thu Jan 29 11:59:48 2015, manasa, Update, General, Beat note frequency discrepancy Freq_compare1.pngFreq_compare2.png
Message ID: 10918     Entry time: Sat Jan 17 15:07:28 2015     Reply to this: 10919   10956
Author: manasa 
Type: Update 
Category: General 
Subject: Beat note frequency discrepancy 

I was around the PSL table and the X end table today.

X end table:

I was at the X end table making some distance measurements for the telescope to the fiber coupler. I have NOT moved anything as yet.

PSL table:

I wanted to get some data to look at the beat note frequency discrepancy between the green and IR. 

I tried making measurements using the spectrum analyzer at the PSL table but the MC was getting unlocked quite often with PSL enclosure open and HEPA on high. 

So I switched off the power supplies to the RFPDs (3 of them) on the PSL table and disconnected the Xmon input to the adder (at the IOO rack) which brings the green beat note signals to the conrol room. I connected the fiber RFPD output to the adder and took a bunch of measurements of the green and IR beat notes from the spectrum analyzer in the control room. I have NOT undone this setup assuming there are no locking plans for tonight and I will come back tomorrow. 

If anyone is planning to do some locking in the meantime, you can undo the connections keeping in mind to turn off the power to the RFPDs before you do so.

P.S. I walked in this morning to PSL FSS slow actuator at 0.89 and brought it down to zero before making measurements.
I also touched the steering mirrors that are part of the Y green beat note alignment while looking at the amplitude of the RF signal on the spectrum analyzer.

ELOG V3.1.3-